# Bridging Quantum Mechanics to Organic Liquid Properties via a Universal Force Field

### v3

@@ Abstract
Molecular dynamics (MD) simulations have emerged as one of the most important cornerstones of modern materials and biological research, offering an indispensable tool for probing condensed-phase systems at the atomic level [29, 20, 53]. By revealing microscopic details of complex phenomena such as structural rearrangements and molecular diffusion—often at resolutions inaccessible to experiments—MD simulations provide critical insights for applications ranging from drug discovery to material engineering [41, 47, 1]. Among these applications, MD simulations play a vital role in the screening and design of organic molecules, such as electrolytes for novel batteries [16, 17] and deep eutectic solvents for green chemistry [50]. They enable the rapid exploration of vast chemical and formula spaces for organic materials, which is impossible to enumerate through experiments.
@@
The accuracy of these simulations, however, is critically dependent on the force field—a mathematical model describing interatomic interactions. This dependency necessitates the development of universal force fields combining accuracy and efficiency. Traditional force fields, such as Amber [22], CHARMM [7], and OPLS [26, 25], model interatomic interactions using a set of simple and predefined functional forms and tabulated parameters. Conventional wisdom indicates that the accuracy of traditional force fields for macroscopic properties essentially relies on error cancellation at microscopic level, a consequence of their limited expressive power. To achieve this cancellation, the force field parameters derived from low-level quantum mechanics (QM) calculations need to be further optimized by experimental data, including spectroscopic measurements [19], and thermodynamic properties such as densities and evaporation enthalpies [26, 25].
@@
In contrast, machine learning (ML)-based force fields utilize the powerful fitting capabilities of neural networks to directly learn interatomic interactions from QM calculations, thereby achieving high accuracy in predicting intramolecular energy landscapes, spectroscopic properties, and related phenomena [31, 32, 27, 39]. However, comparing to traditional force fields, ML force fields without physical constraints require tremendous QM data for training. The development of a universal many-body ML force field is exceptionally challenging due to the vast chemical space of organic molecules and the daunting computational cost of generating sufficient ab initio data. Consequently, prominent ML force fields like MACE-OFF can suffer from accuracy and transferability issues, resulting in inferior performance in bulk property predictions when compared to well-established traditional ones [32]. While other efforts, such as the BAMBOO force field  [17], have demonstrated state-of-the-art (SOTA) performance in predicting the bulk properties of electrolytes, they still depend on experimental density data for empirical fine-tuning [17]. Ultimately, the predictive capability of both traditional and ML force fields remains constrained by the availability and integration of experimental information.
@@
To address this challenge, efforts have been made toward establishing ab initio force fields by refining traditional functional forms to better incorporate the underlying physics of molecular interactions [52, 23]. These force fields are typically based on separating the intermolecular interactions into distinct components [43, liu2019amoeba+] or employing many-body expansion [2]. Despite the solid physical foundation, such force fields often struggle with computational efficiency and transferability [52]. Polarizable force fields, including AMOEBA [40, 9, liu2019amoeba+], APPLE&P [4] and CL&Pol [16], have emerged as a successful framework that retains computational efficiency while capturing the response of electron density to different electrical environments—a critical feature for systems like electrolytes [30]. Nevertheless, parameterizing these polarizable force fields remains a complex, labor-intensive process that often relies on manual adjustments. As a result, their transferability across different chemical environments is uncertain, failing to consistently outperform traditional force fields when applied to systems with varying electronic structures or bonding characteristics.
@@
In summary, currently there is still an urgent need for an universal organic force field that possesses both the accuracy of high-level ab initio calculations, as well as running speed, generalizability, and data efficiency required by large-scale simulations of complex materials. In response to this challenge, we introduce ByteFF-Pol, a graph neural network [10] (GNN)-parameterized polarizable force field that represents a significant advancement in force field development. Unlike traditional and most ML-based force fields, ByteFF-Pol is trained exclusively on high-level QM labels, eliminating the need for experimental calibration. Through a novel force field architecture and training scheme, ByteFF-Pol achieves exceptional accuracy in predicting the thermodynamic and transport properties of small-molecule liquids and electrolytes, outperforming SOTA traditonal and ML force fields. Its zero-shot prediction capabilities enable the exploration of previously intractable chemical spaces, opening up new avenues for data-driven discovery of materials with optimized properties for applications in electrolyte design, complex solvents, and beyond.
@@ 1.1 Overview of ByteFF-Pol
The framework of ByteFF-Pol is illustrated in Figure 1. Unlike traditional methods that rely on look-up tables, ByteFF-Pol employs a GNN model to predict force field parameters directly from molecular graphs. During the training stage, these parameters are integrated with atomic coordinates into energy functions to compute the decomposed energy terms, which are then fitted to reference labels from the absolutely localized molecular orbital energy decomposition analysis (ALMO-EDA) calculations [28]. After training, the parameters of the GNN model are fixed. The force field parameters predicted by the GNN model can be directly used by standard MD engines (e.g., OpenMM [13]) to execute MD simulations. This approach streamlines the parameterization workflow and enhances the transferability of the force field across diverse chemical environments.
@@
Similar to ByteFF [55], the GNN model is composed of three layers. First, a feature layer extracts information about atoms and bonds from molecular graphs to construct atom and bond embeddings ($x_{n}$ and $x_{e}$). Next, these embeddings are propagated through a multi-layer edge-augmented graph transformer (EGT)[24] to produce the hidden representations of atoms and bonds ($h_{n}$ and $h_{e}$), describing local chemical environments. Finally, a pooling layer processes these hidden representations to generate bonded and non-bonded force field parameters. The GNN model carefully considers molecular symmetries in its 2D topology, so the predicted force field parameters maintain these symmetries.
@@
The bonded energy term, $U^{\mathrm{FF}}_{\mathrm{bonded}}$, is consistent with the functional forms used in ByteFF and GAFF2, encompassing standard bond, angle, proper dihedral, and improper dihedral potentials.
The non-bonded energy, $U^{\mathrm{FF}}_{\mathrm{non-bonded}}$, consists of five components: repulsion ($U^{\mathrm{FF}}_{\mathrm{rep}}$), dispersion ($U^{\mathrm{FF}}_{\mathrm{disp}}$), permanent electrostatic ($U^{\mathrm{FF}}_{\mathrm{est}}$), polarization ($U^{\mathrm{FF}}_{\mathrm{pol}}$), and charge transfer ($U^{\mathrm{FF}}_{\mathrm{ct}}$) terms:
@@ 1.2 Fitting to ALMO-EDA labels
Figure 2: ByteFF-Pol is validated against DFT references. (a) Comparison of interaction energies predicted by ByteFF-Pol versus DFT on the validation set (a random sampled subset of 76 k conformations). (b, c) Dimer scan of (b) two EMC molecules and (c) an EMC molecule with a lithium ion (solid lines for ByteFF-Pol and dashed lines for DFT). (d) Cluster of a lithium ion surrounded by three DMC molecules and a FSI- anion. The conformations were first relaxed by DFT (light blue carbon) and then by ByteFF-Pol (grey carbon). (e, f) Decomposed binding energies of clusters with neutral molecules only (e) and clusters with neutral molecules and ions (f). The energy terms are accumulated so that the final point equals to the overall binding energy (solid lines for ByteFF-Pol and dashed lines for DFT). The corresponding components and conformations of the clusters are provided in Supplementary Section 2. The full names and SMILES strings corresponding to the molecular abbreviations are listed in Supplementary Table 1.
While ByteFF-Pol is trained exclusively on dimer data, its physically-motivated functional forms and decomposed training strategy make it transferable to many-body clusters. This transferability is demonstrated in Figure 2(d), which shows a lithium ion and its first solvation shell extracted from an MD simulation. The cluster’s geometries were first optimized with DFT at the B3LYP-D3(BJ)/def2-SVPD level [44], represented by the light blue carbon atoms. The conformations were then re-optimized using ByteFF-Pol, represented by the grey carbon atoms. During the force field optimization, the atomic positions were restrained with a weak harmonic force (1 kcal/mol/Å2) to ensure the system remained in the same local minimum. The resulting structures from both methods show excellent alignment, demonstrating the capability of ByteFF-Pol to accurately predict equilibrium geometries for complex clusters.
@@ 1.3 Thermodynamic Properties of Molecular Liquids
The thermodynamic properties of molecular liquids, primarily governed by intermolecular interactions, serve as critical benchmarks for evaluating force field validity. Using ByteFF-Pol, we computed the densities and evaporation enthalpies for a diverse set of nearly one hundred pure molecular liquids, covering a wide range of functional groups to represent different interaction types. The performance of ByteFF-Pol is compared against results from OPLS-AA [26, 25], OpenFF (i.g. OpenFF-unconstrained-2.2.1) [3], and MACE-OFF (i.g., MACE-OFF23(M)) [32], as shown in Figure 3(a) and (b).
@@
For liquid densities, ByteFF-Pol provides accurate predictions comparable to established force fields like OPLS-AA and OpenFF. The minor disagreement with experimental values is mainly due to a slight, systematic overestimation. This overestimation is consistent with the expected impact of neglecting the nuclear quantum effects (NQE), which typically decreases the density of hydrogen-rich molecules by about 2% [33]. We address the NQE by performing Path-Integral Molecular Dynamics (PIMD) simulations [49], with detailed in Supplementary Section 8.6. As shown in Figure 3(a), ByteFF-Pol-NQE has effectively corrected the systematic overestimation of ByteFF-Pol, resulting lower mean absolute error (MAE) and mean absolute percentage error (MAPE) than other force fields.
@@
Predicting the evaporation enthalpy ($\Delta H_{\text{vap}}$), however, is more challenging due to experimental uncertainties and the ideal gas approximation used for the gas phase simulations (see Supplementary Section 8.2). Despite these difficulties, ByteFF-Pol demonstrates remarkable accuracy, outperforming both OPLS-AA and OpenFF. With the inclusion of NQE, ByteFF-Pol-NQE yields even better results. This high accuracy is particularly noteworthy given that ByteFF-Pol is trained exclusively on DFT dimer data, whereas OPLS-AA and OpenFF are parameterized using experimental data including densities and evaporation enthalpies.
@@
When compared to MACE-OFF, a SOTA ML force field trained purely on DFT data, ByteFF-Pol exhibits substantially lower prediction errors for both density and evaporation enthalpy. This discrepancy highlights the effectiveness of the physically-motivated model and training strategy employed by ByteFF-Pol. Additionally, due to its relatively simple functional form, ByteFF-Pol achieves a simulation speed of 50 ns/day for a 10,000-atom system on an L20 GPU, which is far more efficient than typical ML force fields (several ns/day) [32, 17].
@@
Figure 3: Benchmarks on the thermodynamic properties of molecular liquids. Comparison of different force fields on (a) density and (b) evaporation enthalpy of pure molecular liquids from QUBEKit [23]; (c) density of binary molecular liquids from ThermoML [8]; and (d) density prediction by ByteFF-Pol on additional pure molecular liquids from CRC handbook [35]. The results of MACE-OFF were acquired from Ref 32. Pearson correlation coefficients (r) of each force field are provided in the legends.
@@
Density (g/mL) from QUBEKit [23] $\Delta H_{\mathrm{vap}}$ (kcal/mol) from QUBEKit [23]
@@
OPLS-AA 96 0.024 2.5% 0.99 96 1.64 16.0% 0.85
OpenFF 96 0.024 2.3% 0.99 96 1.30 12.8% 0.87
MACE-OFF 93 0.086 9.4% 0.87 81 2.15 23.2% 0.86
ByteFF-Pol 96 0.029 3.0% 0.98 96 1.13 11.2% 0.90
@@
Density (g/mL) from ThermoML [8] Density (g/mL) from CRC handbook [35]
@@
OPLS-AA 594 0.018 2.02% 0.98 - - - -
OpenFF 592 0.034 3.78% 0.95 - - - -
ByteFF-Pol 595 0.028 3.14% 0.99 2152 0.032 3.28% 0.98
The accuracy of ByteFF-Pol is further validated by its density predictions for binary molecular mixtures, as shown in Figure 3(c). In these systems, ByteFF-Pol demonstrates excellent accuracy, comparable to that of OPLS-AA and OpenFF. A slight, systematic overestimation of density across most mixtures can again be attributed to the neglect of NQE. However, due to the substantial computational cost of PIMD simulations, we did not perform PIMD simulations for this or subsequent benchmarks. Minor underestimations are observed in some high-density systems containing polyhalogenated molecules, which is likely due to the inherent limitations of the model’s functional form in capturing anisotropic interactions such as halogen bonds and $\sigma$-holes. Nevertheless, ByteFF-Pol demonstrates reliable overall performance, effectively modeling intermolecular forces across diverse chemical environments.
The transferability and universality of ByteFF-Pol are evaluated using density data for pure molecular liquids sourced from CRC handbook (Figure 3 (d)). Despite being trained on fewer than 400 molecules, ByteFF-Pol delivers precise density predictions for over 2,000 distinct molecules. The data points are colored based on the similarity of the local atomic chemical environments between the tested molecules and the training set, gauged by the similarity of the embedding vectors. As shown in Figure 3 (d), the densities of most molecules with large similarity to the training set can be predicted accurately by ByteFF-Pol. This observation underscores that ByteFF-Pol’s transferability stems from its ability to capture the chemical environments of atoms, rather than explicit training on entire molecular structures. This demonstrates that comprehensive coverage of atomic chemical environments in the training set is sufficient to enable accurate predictions across chemically diverse molecules. Overall, ByteFF-Pol demonstrates exceptional performance in predicting the thermodynamic properties of molecular liquids with promising transferability.

### v4

@@ Abstract
Molecular dynamics (MD) simulations have emerged as one of the most important cornerstones of modern materials and biological research, offering an indispensable tool for probing condensed-phase systems at the atomic level [kholmurodov2007molecular, hollingsworth2018molecular, yan2023molecular]. By revealing microscopic details of complex phenomena such as structural rearrangements and molecular diffusion—often at resolutions inaccessible to experiments—MD simulations provide critical insights for applications ranging from drug discovery to material engineering [reymond2012exploring, tafrishi2022molecular, alonso2006combining]. Among these applications, MD simulations play a vital role in the screening and design of organic molecules, such as electrolytes for novel batteries [golovizninaExtensionCLPolPolarizable2021, gong2025predictive] and deep eutectic solvents for green chemistry [https://doi.org/10.1002/wcms.1598]. They enable the rapid exploration of vast chemical and formula spaces for organic materials, which is impossible to enumerate through experiments.
@@
The accuracy of these simulations, however, is critically dependent on the force field—a mathematical model describing interatomic interactions. This dependency necessitates the development of universal force fields combining accuracy and efficiency. Traditional force fields, such as Amber [hornak2006comparison], CHARMM [brooks2009charmm], and OPLS [jorgensen1984optimized, jorgensen1986optimized], model interatomic interactions using a set of simple and predefined functional forms and tabulated parameters. Conventional wisdom indicates that the accuracy of traditional force fields for macroscopic properties essentially relies on error cancellation at microscopic level, a consequence of their limited expressive power. To achieve this cancellation, the force field parameters derived from low-level quantum mechanics (QM) calculations need to be further optimized by experimental data, including spectroscopic measurements [han2025refining], and thermodynamic properties such as densities and evaporation enthalpies [jorgensen1984optimized, jorgensen1986optimized].
@@
In contrast, machine learning (ML)-based force fields utilize the powerful fitting capabilities of neural networks to directly learn interatomic interactions from QM calculations, thereby achieving high accuracy in predicting intramolecular energy landscapes, spectroscopic properties, and related phenomena [kovacsEvaluationMACEForce2023, kovacsMaceOff2025, juSevenNet2025, pleFoundationModelAccurate2025]. However, comparing to traditional force fields, ML force fields without physical constraints require tremendous QM data for training. The development of a universal many-body ML force field is exceptionally challenging due to the vast chemical space of organic molecules and the daunting computational cost of generating sufficient ab initio data. Consequently, prominent ML force fields like MACE-OFF can suffer from accuracy and transferability issues, resulting in inferior performance in bulk property predictions when compared to well-established traditional ones [kovacsMaceOff2025]. While other efforts, such as the BAMBOO force field [gong2025predictive], have demonstrated state-of-the-art (SOTA) performance in predicting the bulk properties of electrolytes, they still depend on experimental density data for empirical fine-tuning [gong2025predictive]. Ultimately, the predictive capability of both traditional and ML force fields remains constrained by the availability and integration of experimental information.
@@
To address this challenge, efforts have been made toward establishing ab initio force fields by refining traditional functional forms to better incorporate the underlying physics of molecular interactions [xu2018perspective, hortonQUBEKitAutomatingDerivation2019]. These force fields are typically based on separating the intermolecular interactions into distinct components [schmidt_transferable_2015, liu2019amoeba+] or employing many-body expansion [beran2009approximating]. Despite the solid physical foundation, such force fields often struggle with computational efficiency and transferability [xu2018perspective]. Polarizable force fields, including AMOEBA [ponderCurrentStatusAMOEBA2010, chungClassicalExchangePolarization2022, liu2019amoeba+], APPLE&P [borodinPolarizableForceField2009] and CL&Pol [golovizninaExtensionCLPolPolarizable2021], have emerged as a successful framework that retains computational efficiency while capturing the response of electron density to different electrical environments—a critical feature for systems like electrolytes [kirshCriticalEvaluationPolarizable2024]. Nevertheless, parameterizing these polarizable force fields remains a complex, labor-intensive process that often relies on manual adjustments. As a result, their transferability across different chemical environments is uncertain, failing to consistently outperform traditional force fields when applied to systems with varying electronic structures or bonding characteristics.
@@
In summary, currently there is still an urgent need for a universal organic force field that possesses both the accuracy of high-level ab initio calculations, as well as running speed and generalizability required by large-scale simulations of complex materials. In response to this challenge, we introduce ByteFF-Pol, a graph neural network [corso2024graph] (GNN)-parameterized polarizable force field that represents a significant advancement in force field development. Unlike traditional and most ML-based force fields, ByteFF-Pol is trained exclusively on high-level QM labels, eliminating the need for experimental calibration. Through a novel force field architecture and training scheme, ByteFF-Pol achieves exceptional accuracy in predicting the thermodynamic and transport properties of small-molecule liquids and electrolytes, outperforming SOTA traditional and ML force fields. Its zero-shot prediction capabilities enable the exploration of previously intractable chemical spaces, opening up new avenues for data-driven discovery of materials with optimized properties for applications in electrolyte design, complex solvents, and beyond.
@@ 1.1 Overview of ByteFF-Pol
The ByteFF-Pol framework, depicted in Figure 1, utilizes a Graph Neural Network (GNN) to derive force field parameters directly from molecular graphs. This modern parameterization technique offers significant advantages over traditional look-up tables [wangEndtoendDifferentiableConstruction2022, thurlemann2023regularized, zheng2025data, seute2025grappa]. During training, the predicted force field parameters are integrated with atomic coordinates into energy functions to compute the decomposed energy terms. And the GNN model is optimized by fitting the decomposed energy terms to reference data from the absolutely localized molecular orbital energy decomposition analysis (ALMO-EDA) calculations [khaliulinEDAv1Original2007]. Once trained, the finalized GNN model can generate parameters for direct use in standard molecular dynamics (MD) engines (e.g., OpenMM [eastman2017openmm]) to execute MD simulations. This approach streamlines the parameterization workflow and enhances the transferability of the force field across diverse chemical environments.
@@
Similar to ByteFF [zheng2025data], the GNN model is composed of three layers. First, a feature layer extracts information about atoms and bonds from molecular graphs to construct atom and bond embeddings ($x_{n}$ and $x_{e}$). Next, these embeddings are propagated through a multi-layer edge-augmented graph transformer (EGT)[hussainGlobalSelfAttentionReplacement2022] to produce the hidden representations of atoms and bonds ($h_{n}$ and $h_{e}$), describing local chemical environments. Finally, a pooling layer processes these hidden representations to generate bonded and non-bonded force field parameters. The GNN model carefully considers molecular symmetries in its 2D topology, so the predicted force field parameters maintain these symmetries.
@@
The bonded energy term, $U^{\mathrm{FF}}_{\mathrm{bonded}}$, is consistent with the functional forms used in ByteFF and GAFF2, encompassing standard bond, angle, proper dihedral, and improper dihedral potentials. The non-bonded energy, $U^{\mathrm{FF}}_{\mathrm{non-bonded}}$, consists of five components: repulsion ($U^{\mathrm{FF}}_{\mathrm{rep}}$), dispersion ($U^{\mathrm{FF}}_{\mathrm{disp}}$), permanent electrostatic ($U^{\mathrm{FF}}_{\mathrm{est}}$), polarization ($U^{\mathrm{FF}}_{\mathrm{pol}}$), and charge transfer ($U^{\mathrm{FF}}_{\mathrm{ct}}$) terms:
@@
ByteFF-Pol shares a similar GNN model with our previous work, ByteFF, for predicting force field parameters. The major advancements lie in the parametrization methodology of non-bonded parameters: instead of distilling the conventional GAFF2 force field, we now fit to QM data directly. This constitutes a non-trivial task, as most SOTA force fields require experimental fine-tuning due to their inaccurate function form and incomplete QM dataset. To address these challenges, we curated a large-scale, chemically diverse QM dataset, generating decomposed energy labels using the ALMO-EDA method in addition to the total energy. The EDA data encodes physically meaningful information that substantially enhances the robustness of the resulting model. Correspondingly, we designed EDA-compatible force field functions and a tailored training protocol to ensure both fitting efficiency and accuracy. In the subsequent sections, we detail the training procedure of ByteFF-Pol and demonstrate its exceptional performance in both microscopic energy predictions and bulk simulations through comprehensive benchmarking studies.
@@ 1.2 Fitting to ALMO-EDA labels
Figure 2: ByteFF-Pol is validated against DFT references. (a) Comparison of interaction energies predicted by ByteFF-Pol versus DFT on the validation set (a random sampled subset of 76 k conformations). (b, c) Dimer scan of (b) two EMC molecules and (c) an EMC molecule with a lithium ion (solid lines for ByteFF-Pol and dashed lines for DFT). (d) Cluster of a lithium ion surrounded by three DMC molecules and a FSI- anion. The conformations were first relaxed by DFT (light blue carbon) and then by ByteFF-Pol (grey carbon). (e, f) Decomposed binding energies of clusters with neutral molecules only (e) and clusters with neutral molecules and ions (f). The energy terms are accumulated so that the final point equals to the overall binding energy (solid lines for ByteFF-Pol and dashed lines for DFT). The corresponding components and conformations of the clusters are provided in Supplementary Section 2.1. The full names and SMILES strings corresponding to the molecular abbreviations are listed in Supplementary Table 1.
While ByteFF-Pol is trained exclusively on dimer data, its physically-motivated functional forms and decomposed training strategy make it transferable to many-body clusters. This transferability is demonstrated in Figure 2(d), which shows a lithium ion and its first solvation shell extracted from an MD simulation. The cluster’s geometries were first optimized with DFT at the B3LYP-D3(BJ)/def2-SVPD level [stephens1994ab], represented by the light blue carbon atoms. The conformations were then re-optimized using ByteFF-Pol, represented by the grey carbon atoms. During the force field optimization, the atomic positions were restrained with a weak harmonic force (1 kcal/mol/Å2) to ensure the system remained in the same local minimum. The resulting structures from both methods show excellent alignment, demonstrating the capability of ByteFF-Pol to accurately predict equilibrium geometries for complex clusters.
@@ 1.3 Thermodynamic Properties of Molecular Liquids
The thermodynamic properties of molecular liquids, primarily governed by intermolecular interactions, serve as critical benchmarks for evaluating force field validity. Using ByteFF-Pol, we computed the densities and evaporation enthalpies for a diverse set of nearly one hundred pure molecular liquids, covering a wide range of functional groups to represent different interaction types. The performance of ByteFF-Pol is benchmarked against that of SOTA force fields, including a conventional force field (i.e. OPLS-AA [jorgensen1984optimized, jorgensen1986optimized]), a polarizable force field (i.e. AMOEBA [ponderCurrentStatusAMOEBA2010, walkerAutomationAMOEBAPolarizable2022]), and a ML force field (i.e., MACE-OFF23(M)) [kovacsMaceOff2025], as shown in Figure 3(a) and (b). Parmetrization and simulation details can be found in Supplementary Section 8, and comparsion with additional force fields are provided in Supplementary Section 9.1.
@@
For liquid densities, ByteFF-Pol provides accurate predictions comparable to established force fields like OPLS-AA and AMOEBA. The minor disagreement with experimental values is mainly due to a slight, systematic overestimation. This overestimation is consistent with the expected impact of neglecting the nuclear quantum effects (NQEs), which typically decreases the density of hydrogen-rich molecules by about 2% [kurnikov2024neural]. We address the NQEs by performing Path-Integral Molecular Dynamics (PIMD) simulations [tuckerman_statistical_2023], with detailed in Supplementary Section 8.6. As shown in Figure 3(a), ByteFF-Pol-NQE has effectively corrected the systematic overestimation of ByteFF-Pol, resulting a lower mean absolute error (MAE) and mean absolute percentage error (MAPE) than other force fields.
@@
Predicting the evaporation enthalpy ($\Delta H_{\text{vap}}$), however, is more challenging due to experimental uncertainties and the ideal gas approximation used for the gas phase simulations (see Supplementary Section 8.2). Despite these difficulties, ByteFF-Pol demonstrates remarkable accuracy, outperforming both OPLS-AA and AMOEBA. With the inclusion of NQEs, ByteFF-Pol-NQE yields even better results. This high accuracy is particularly noteworthy given that ByteFF-Pol is trained exclusively on DFT dimer data, whereas OPLS-AA and AMOEBA are parameterized using experimental data including densities and evaporation enthalpies.
@@
When compared to MACE-OFF, a SOTA ML force field trained purely on DFT data, ByteFF-Pol exhibits substantially lower prediction errors for both density and evaporation enthalpy. As shown in Supplementary Figure 14, equipped with physics-driven long-range interactions, MACELES-OFF [kim2025universal] significantly outperforms MACE-OFF, although it still exhibits larger errors than ByteFF-Pol. This discrepancy highlights the effectiveness of the physically-motivated model and training strategy employed by ByteFF-Pol.
@@
Figure 3: Benchmarks on the thermodynamic properties of molecular liquids. Comparison of different force fields on (a) density and (b) evaporation enthalpy of pure molecular liquids from QUBEKit [hortonQUBEKitAutomatingDerivation2019]; (c) density of binary molecular liquids from ThermoML [chirico2003thermoml]; and (d) density prediction by ByteFF-Pol on additional pure molecular liquids from CRC handbook [Lide2004CRC]. The results of MACE-OFF were acquired from Ref kovacsMaceOff2025. Pearson correlation coefficients (r) of each force field are provided in the legends.
@@
Density (g/mL) from QUBEKit [hortonQUBEKitAutomatingDerivation2019] $\Delta H_{\mathrm{vap}}$ (kcal/mol) from QUBEKit [hortonQUBEKitAutomatingDerivation2019]
@@
OPLS-AA 96 0.024 2.5% 0.99 96 1.52 14.8% 0.87
AMOEBA 87 0.025 2.4% 0.99 87 0.85 8.1% 0.91
MACE-OFF 92 0.084 9.8% 0.92 80 2.14 23.1% 0.86
ByteFF-Pol 96 0.029 3.0% 0.98 96 0.95 9.3% 0.94
@@
Density (g/mL) from ThermoML [chirico2003thermoml] Density (g/mL) from CRC handbook [Lide2004CRC]
@@
OPLS-AA 594 0.018 2.0% 0.98 - - - -
AMOEBA 554 0.018 2.0% 0.99 - - - -
ByteFF-Pol 595 0.028 3.1% 0.99 2152 0.032 3.28% 0.98
ByteFF-Pol-NQE 595 0.017 1.8% 0.99 - - - -
The accuracy of ByteFF-Pol is further validated by its density predictions for binary molecular mixtures, as shown in Figure 3(c). In these systems, ByteFF-Pol demonstrates excellent accuracy, comparable to that of OPLS-AA and AMOEBA. A slight, systematic overestimation of density by ByteFF-Pol can again be corrected by incorporating the NQEs. However, due to the substantial computational cost of PIMD simulations, we did not perform PIMD simulations for this or subsequent benchmarks. Minor underestimations are observed in some high-density systems containing polyhalogenated molecules, which is likely due to the inherent limitations of the model’s functional form in capturing anisotropic interactions such as halogen bonds and $\sigma$-holes. Nevertheless, ByteFF-Pol demonstrates reliable overall performance, effectively modeling intermolecular forces across diverse chemical environments.
The transferability and universality of ByteFF-Pol are evaluated using density data for pure molecular liquids sourced from CRC handbook (Figure 3 (d)). Despite being trained on fewer than 400 molecules, ByteFF-Pol delivers precise density predictions for over 2,000 distinct molecules. The data points are colored based on the similarity of the local atomic chemical environments between the tested molecules and the training set, gauged by the similarity of the embedding vectors. As shown in Figure 3 (d), the densities of most molecules with large similarity to the training set can be predicted accurately by ByteFF-Pol. As demonstrated in Supplementary Section 3.2, a clear correlation exists between the similarity score and the density prediction accuracy. This observation underscores that ByteFF-Pol’s transferability stems from its ability to capture the chemical environments of atoms, rather than explicit training on entire molecular structures. This indicates that comprehensive coverage of atomic chemical environments in the training set is sufficient to enable accurate predictions across chemically diverse molecules (see Supplementary Section 10 for a detailed analysis of the training set’s chemical space). Overall, ByteFF-Pol demonstrates exceptional performance in predicting the thermodynamic properties of molecular liquids with promising transferability.